# The continuum limit of some products of random matrices associated with renewing flows

### v1

@@ Abstract.
We consider the continuum limit of some products of random matrices in $\text{SL}(d,{\mathbb{R}})$ that arise as discretisations of incompressible renewing flows— that is, of flows corresponding to a divergence-free velocity field that takes independent, identically-distributed values in successive time intervals of duration proportional to $\tau$. The statistical properties of the product are encoded in its generalised Lyapunov exponent whose computation reduces to finding the leading eigenvalue of a certain transfer operator. In the continuum (Kraichnan–Kazantsev) limit obtained by neglecting the terms of order $o(\tau^{2})$, the transfer operator becomes a partial differential operator and, for a certain type of disorder which we call “symmetric”, some calculations are feasible. For $d=2$, we compute the growth rate of the product in terms of complete elliptic integrals. By letting the elliptic modulus vary, we obtain a spectral problem, corresponding to a modulus-dependent random renewing flow, which may be viewed as a perturbation of the spectral problem for the angular Laplacian. In this way, we deduce expansions for the generalised Lyapunov exponent in ascending powers of the modulus. These expansions generalise to the case $d\geq 2$, and we compute the first few terms explicitly for $d\in\{2,3\}$.
@@
This work was initiated during a visit to the Laboratoire de Physique Théorique et Modèles Statistiques in the summer of 2023. I am grateful to the Université Paris–Saclay for supporting the visit financially, and to my long-time collaborators Alain Comtet and Christophe Texier for many helpful discussions.
@@ 1. Introduction
Products of random matrices arise as mathematical models of many disordered physical systems [4]. However, the fact that the elements in such products do not commute makes the study of their statistical properties challenging. In a recent publication [2], we exhibited one case where detailed calculations were possible— namely a product of random matrices in $\text{SL}(2,{\mathbb{R}})$ that models a quantum particle in a one-dimensional potential with randomly distributed impurities; progress was made by taking the continuum limit and exploiting the specific features of the disorder. The present paper pursues a similar objective; only, the product under scrutiny concerns toy models for turbulent incompressible fluid flow in $d$-dimensional space, so that the matrices in the product belong to the group $\text{SL}(d,{\mathbb{R}})$. Roughly speaking, the special features that enable calculations are as follows: (1) we consider a continuum limit of the product; (2) we assume that the disorder takes a highly symmetric form; (3) we embed the problem in a parameter-dependent family that contains a trivially solvable case.
@@ 1.1. Renewing flows
We consider a $d$-dimensional random velocity field with components of the form
@@
where the $u_{ij}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ are periodic functions with a common period, zero mean over a period, and the $\{\eta_{ij}(t)\}_{t\geq 0}$ are independent copies of the same random process $\{\eta(t)\}_{t\geq 0}$. The corresponding flow, governed by
(1.2) $\dot{x}={u}(x,t)\,,\;\;t>0\,$,
has been used in simplified models for the diffusion of a scalar (say a temperature or the density of some pollutant) carried by a “turbulent” flow which is assumed to be random but unaffected by the presence of the scalar. The simplest examples occur when the $u_{ij}$ are expressible in terms of trigonometric polynomials [3, 15]. If we consider the paths followed by particles near a given trajectory $x=x(t)$ , we see that the separation, say $\delta x(t)$, between the paths evolves locally according to the equations
$\frac{d}{dt}\delta x_{i}=\sum_{j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\delta x_{j}\,,\;\;1\leq i\leq d\,.$
One function of interest is then
(1.3) ${\mathbb{R}}\ni\ell\mapsto\lim_{t\rightarrow\infty}\frac{\ln{\mathbb{E}}\left[\left|\frac{\delta x(t)}{\delta x(0)}\right|^{\ell}\right]}{t}$
where $|\cdot|$ denotes the usual Euclidean norm in ${\mathbb{R}}^{d}$. Its Legendre transform is the rate function that describes the large-deviation statistics of $\delta x(t)$. The question addressed by Haynes & Vanneste [10] is: “Does knowledge of this function help in characterising the long-time behaviour of the passive scalar?” For a diffusion process subject to periodic boundary conditions, they showed that the answer is “yes”, provided that the period is a small enough multiple of that of the velocity field.
In this paper, we shall not address the diffusion problem but instead concentrate on the dynamical system (1.2). Let us assume that, after selecting a reference length, we measure length in dimensionless units, so that the resulting fluid model has the following properties:
@@
    The spatial period of the $u_{ij}$ is $2\pi$.
@@
    For every $t\geq 0$, $\eta(t)$ is uniformly-distributed in $(-\pi,\pi)$.
@@
$t_{n+1}-i\,\delta t<t<t_{n+1}-(i-1)\,\delta t\,,\;\;i\in\{1,\,2,\,\ldots\,,d\}\,$,
the only non-vanishing velocity component is the $i$th one. Such flows are called “random flows with renewal” or, simply, “renewing flows” [18].
@@ 1.2. Products of matrices from renewing flows
A particular feature of the products we consider is the fact, proved in Appendix A, that the $\xi_{ij}$ are independent. We are thus led to studying certain products
@@
of random independent and identically-distributed matrices with a particular structure: each $g_{n}$ is itself a product of random independent triangular matrices of unit determinant and hence belongs to the Lie group $\text{SL}(d,{\mathbb{R}})$— a consequence of the fact that the velocity field defined by Equation (1.1) is divergence-free.
@@ 1.3. The generalised Lyapunov exponent
Associated with the product (1.10) is the counterpart of the quantity (1.3):
(1.11) $L(\ell):=\lim_{n\rightarrow\infty}\frac{1}{n}\ln{\mathbb{E}}\left[1_{\ell}\left(x\Pi_{n}\right)\right]=\sum_{j=1}^{\infty}\gamma_{j}\frac{\ell^{j}}{j!}$
@@
$L$ may be viewed as a cumulant generating function associated with the dimensionless random variable $\ln\left|{x}\Pi_{n}\right|$; it is in fact independent of the x{x} appearing in its definition. In terms of the renewing flow, $\gamma_{j}/(\delta t\,d)$ is the discrete counterpart of
@@
so that, in particular, $\gamma_{1}/(\delta t\,d)$ measures the rate at which nearby trajectories separate. More generally, $L$ encodes the large-$n$ statistical properties of the product: if $\text{\rm d}\ell$ denotes an interval of infinitesimal length $d\ell$ centered at $\ell\in{\mathbb{R}}$, then
(1.13) ${\mathbb{P}}\left(\frac{1}{n}\ln\left|x\Pi_{n}\right|\in\text{\rm d}\ell\right)\propto\exp\left[-n\,L^{\ast}(\ell)\right]\,d\ell\;\;\text{as $n\rightarrow\infty$}\,.$
where
(1.14) $L^{\ast}(\ell):=\min_{p}\left[p\ell-L(p)\right]$
is the Legendre transform of $L$.
@@ 1.4. The transfer operator
where $\lambda(\ell)$ is the dominant eigenvalue of the transfer operator.
@@ 1.5. The continuum limit
Associated with a given random velocity field is the matrix-valued correlation function with entries
${\mathbb{E}}\left[u_{i}(x,t)\,u_{j}(x^{\prime},t^{\prime})\right]\,,\;\;1\leq i,j\leq d\,.$
For the piecewise-constant-in-time velocity field considered in this paper, only the diagonal entries are non-zero; a simple calculation shows that they take the form
(1.19) ${\mathbb{E}}\left[u_{i}(x,t)\,u_{i}(x^{\prime},t^{\prime})\right]=\sum_{n=0}^{\infty}{\mathbf{1}}_{I_{n,i}}(t)\,{\mathbf{1}}_{I_{n,i}}(t^{\prime})\sum_{j\neq i}b_{ij}(x_{j}-x_{j}^{\prime})\,,\;\;1\leq i\leq d\,$,
where the $b_{ij}$ are cosine series with positive coefficients, and ${\mathbf{1}}_{I_{n,i}}$ is the characteristic function of the interval $I_{n,i}:=\left(t_{n+1}-i\,\delta t,\,t_{n+1}-(i-1)\,\delta t\right)$. If we multiply this identity by a suitable test function of $t^{\prime}$ and integrate over $t^{\prime}$, we see that the limit $\delta t\rightarrow 0+$, henceforth referred to as the continuum limit— or Kraichnan–Kazantsev limit, following the terminology used in the fluid dynamics literature— corresponds to a velocity field that is delta-correlated in time.
@@ 1.6. Outline of the paper and main results
For small $\tau$, Equation (1.18) then implies that
(1.26) $L(\ell)=\sum_{j=1}^{\infty}\gamma_{j}\frac{\ell^{j}}{j!}=L\left(k,\ell\right){\Bigl|_{k=1/\sqrt{d}}}$
@@
(1.27) $L(k,\ell):=\tau^{2}\left[(d-1)\,\ell+\frac{d-1}{d}\,\ell^{2}+\mu(k,\ell)\right]$
and $\mu(k,\ell)$ is the particular eigenvalue that maximises the expression on the right-hand side. We take advantage of the fact that the companion spectral problem (1.25) is completely solvable when $k=0$ by embedding our original problem in a parameter-dependent family, where the parameter $k$ assumes “small” values. To facilitate the calculations, we work with the Iwasawa realisation, described in §3, of the basic representation, because it is in this realisation that the eigenfunctions of the Laplacian assume their simplest form.
In §4, we study the $k$-dependent spectral problem (1.25) when $d=2$. First, for fixed $0\leq k<1$, we consider its small-$\ell$ limit, and find explicit formulae for the first two cumulants:
(1.28) $\frac{1}{\tau^{2}}\,\gamma_{1}(k)=2\,\frac{\boldsymbol{\mathsf{E}}}{\boldsymbol{\mathsf{K}}}-1$
@@
(1.29) $\frac{1}{\tau^{2}}\,\frac{\gamma_{2}(k)}{2}=\frac{3}{2}-k^{2}-\frac{\boldsymbol{\mathsf{E}}}{\boldsymbol{\mathsf{K}}}+\frac{2\pi^{2}}{\boldsymbol{\mathsf{K}}^{2}}\sum_{j=1}^{\infty}\left(\frac{q^{2j}}{1-q^{4j}}\right)^{2}-\frac{6\pi^{2}}{\boldsymbol{\mathsf{K}}^{2}}\sum_{j=1}^{\infty}\left(\frac{q^{2j-1}}{1-q^{4j-2}}\right)^{2}$
where $\boldsymbol{\mathsf{K}}$ and $\boldsymbol{\mathsf{E}}$ are the complete elliptic integrals of the first and second kind respectively, evaluated at $k$, and $q$ is the corresponding nome. The fact that the parameter $k$ appears in these formulae as the modulus of the elliptic integrals gives some confidence that its introduction may indeed be the “right” way of relating our spectral problem to that for $\Delta_{K}$. Plots of
(1.30) $\ell\mapsto\ell\gamma_{1}+\ell^{2}\frac{\gamma_{2}}{2}=L(\ell)+o\left(\ell^{2}\right)$
@@
(1.31) $\ell\mapsto\frac{(\ell-\gamma_{1})^{2}}{2\gamma_{2}}$
against $\ell$, for $\tau=1$, are shown in Figure 1 (a). In these formulae, which express the fact that the distribution of the random variable $\left|x\Pi_{n}\right|$ is approximately lognormal for large $n$ [4], the $\gamma_{j}$ are evaluated at $k=1/\sqrt{2}$.
(a)(b)
Figure 1. Plots of $L$ (solid blue curve) and $L^{\ast}$ (solid red curve) against $\ell$ for $\tau=1$ and (a) $d=2$ and (b) $d=3$. In (a) are also displayed the approximation (1.30) of $L$ (dashed curve) and the approximation (1.31) of $L^{\ast}$ (dot-dashed curve) which use only the first two cumulants.
Then, by considering the small-$k^{2}$ limit of the spectral problem (1.25), we obtain expansions for $L$ and $L^{\ast}$ in ascending powers of $k^{2}$:
(1.32) $\frac{1}{\tau^{2}}L(k,\ell)=\frac{\ell\left(\ell+2\right)}{2}-\frac{\ell\left(\ell+2\right)}{2}k^{2}\\ +\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{128}k^{4}+\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{256}k^{6}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(7\ell^{4}+28\ell^{3}+36\ell^{2}+16\ell-5248\right)}{2097152}k^{8}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(21\ell^{4}+84\ell^{3}+108\ell^{2}+48\ell-7552\right)}{4194304}k^{10}+\cdots$
@@
(1.33) $\frac{1}{\tau^{2}}L^{\ast}(k,\ell)=\frac{\left(\ell^{\prime}-1\right)^{2}}{2}+\frac{\left(\ell^{\prime}-1\right)\left(\ell^{\prime}+1\right)}{2}k^{2}\\ -\frac{(\ell^{\prime})^{4}-74(\ell^{\prime})^{2}+9}{128}k^{4}-\frac{\left[3(\ell^{\prime})^{2}-14\ell^{\prime}+3\right]\left[3(\ell^{\prime})^{2}+14\ell^{\prime}+3\right]}{256}k^{6}\\ +\frac{7(\ell^{\prime})^{8}+948(\ell^{\prime})^{6}-211974(\ell^{\prime})^{4}+1781972(\ell^{\prime})^{2}-47241}{2097152}k^{8}\\ +\frac{133(\ell^{\prime})^{8}+15244(\ell^{\prime})^{6}-990274(\ell^{\prime})^{4}+4389324(\ell^{\prime})^{2}-67995}{4194304}k^{10}+\cdots$
where $\ell^{\prime}:=\ell/\tau^{2}$ and we have used the following elementary property of the Legendre transform:
$\forall\,\alpha>0\,,\;\;L(\ell)\mapsto\alpha\,L(\ell)\iff L^{\ast}(\ell)\mapsto\alpha\,L^{\ast}\left(\ell/\alpha\right)\,.$
In §5, we turn to the case $d=3$. We do not have the analytical tools to make progress in the small-$\ell$ limit. In the small-$k^{2}$ limit, however, we can once again compute expansions for $L$ and its Legendre transform; we obtain
(1.34) $\frac{1}{\tau^{2}}L(k,\ell)=\frac{2\ell\left(\ell+3\right)}{3}-\frac{4\ell\left(\ell+3\right)}{5}k^{2}\\ +\frac{12\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)}{875}k^{4}\\ -\frac{72\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)\left(9\ell^{2}+27\ell-395\right)}{3128125}k^{6}+\cdots$

### v2

@@ Abstract.
We consider the continuum limit of some products of random matrices in $\text{SL}(d,{\mathbb{R}})$ that arise as discretisations of incompressible renewing flows— that is, of flows corresponding to a divergence-free velocity field that takes independent, identically-distributed values in successive time intervals of duration proportional to $\tau$. The statistical properties of the product are encoded in its generalised Lyapunov exponent whose computation reduces to finding the leading eigenvalue of a certain transfer operator. In the continuum limit obtained by neglecting the terms of order $o(\tau^{2})$, the transfer operator becomes a partial differential operator and, for a certain type of disorder which we call “symmetric”, some calculations are feasible. For $d=2$, we compute the growth rate of the product in terms of complete elliptic integrals. By letting the elliptic modulus vary, we obtain a spectral problem, corresponding to a modulus-dependent random renewing flow, which may be viewed as a perturbation of the spectral problem for the angular Laplacian. In this way, we deduce expansions for the generalised Lyapunov exponent in ascending powers of the modulus. These expansions generalise to the case $d\geq 2$, and we compute the first few terms explicitly for $d\in\{2,3\}$.
@@
ORCID: 0000-0002-9646-0401. This work was initiated during a visit to the Laboratoire de Physique Théorique et Modèles Statistiques in the summer of 2023. I am grateful to the Université Paris–Saclay for supporting the visit financially, and to my long-time collaborators Alain Comtet and Christophe Texier for many helpful discussions. The author has no relevant financial or non-financial interests to disclose. There is no data to be made available.
@@ 1. Introduction
Products of random matrices arise as mathematical models of many disordered physical systems; see Part II of the monograph [8]— and the extensive bibliography therein— where concrete applications to the study of Ising spin chains, transport in random media, chaotic dynamics etc. are described. It is typical of such products that they grow at a deterministic, almost-sure rate as the number of elements in the product increases. This growth rate, known as the Lyapunov exponent of the product, is the most basic quantity of interest in the applications, as it usually provides a quantitative measure of some important aspect of the physical phenomenon being modelled; for instance the reciprocal of the localisation length in systems exhibiting Anderson localisation, or the rate at which nearby trajectories separate in a chaotic system. However, the fact that the elements seldom commute makes its calculation challenging. A brief survey of the exact results that have been obtained in this direction may be found in [6]; such cases are few, and the calculation relies on special features of the disorder— as in Dyson’s random chain model [11] or Lloyd’s tight-binding model [19]— and/or on taking the continuum limit of the product— as in Halperin’s study of a quantum particle in a white noise potential [15]. The continuum limit of the two-by-two case is exceptional: the problem of computing the Lyapunov exponent was effectively solved in [4].
If one looks beyond the almost-sure growth rate and seeks to ascertain the nature of the product’s fluctuations then the relevant quantity is the generalised Lyapunov exponent [8], which will be defined shortly, and whose calculation presents an even greater challenge. In [6], we exhibited one case where detailed calculations were possible, namely the continuum limit of a product of random matrices in the group $\text{SL}(2,{\mathbb{R}})$ that models a quantum particle in a one-dimensional potential with impurities. It is the purpose of the present paper to describe another such case; only, this time, the product under scrutiny concerns toy models for turbulent incompressible fluid flow in $d$-dimensional space, so that the matrices in the product belong to the group $\text{SL}(d,{\mathbb{R}})$. As we shall explain in the remainder of this introduction, the new case features the ingredients necessary to make progress: (1) we consider a continuum limit of the product; (2) we assume that the disorder takes a highly symmetric form; (3) we embed the problem in a parameter-dependent family that contains a trivially solvable case.
1.1. Motivation
@@
The diffusion of a scalar carried along by a flow in $d$-dimensional space is governed by the equation
(1.1) $\frac{\partial\rho}{\partial t}=u\cdot\nabla\rho+D\,\Delta\rho\,.$
Here, $\rho=\rho(x,t)$ denotes, say, a temperature or the density of some pollutant, $D$ is the diffusivity of the fluid in which the scalar is advected, and $u=u(x,t)$ is the fluid’s velocity field. It is assumed that the scalar is “passive”, meaning that its presence in the fluid does not affect the velocity field.
To any given velocity field, one may associate the dynamical system
(1.2) $\dot{x}={u}(x,t)\,,\;\;t>0\,.$
If we consider the paths followed by particles near a given trajectory $x=x(t)$ , we see that the separation, say $\delta x(t)$, between the paths evolves locally according to the equations
(1.3) $\frac{d}{dt}\delta x_{i}=\sum_{j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\delta x_{j}\,,\;\;1\leq i\leq d\,.$
For the case where the velocity field is random, Haynes & Vanneste examine the relationship between the large-deviation statistics of $\delta x(t)$ and the long-time behaviour of the two-point covariance function
(1.4) $C(x,x^{\prime},t):={\mathbb{E}}\left[\rho(x,t)\rho(x^{\prime},t)\right]\,.$
More precisely, they consider divergence-free random velocity fields that are spatially homogeneous, so that the covariance function takes the form
(1.5) $C(x,x^{\prime},t)=\Gamma(x^{\prime}-x,t)$
for some function $\Gamma$ of space and time; their aim is then to ascertain under what conditions knowledge of the flow’s stretching characteristics suffices to determine the decay rate of $\Gamma(0,t)$ as $t\rightarrow\infty$. To this end, they look for velocity fields such that some closed equation for the temporal evolution of $\Gamma(x,t)$ may be derived and analysed. To construct velocity fields that meet these requirements, it is convenient to assume that the scalar evolves in a spatial domain obtained by extending periodically a basic hypercube aligned with the axes, denoting by $P$ the (common) period in each of the coordinates.
1.2. Kraichnan’s velocity field ensemble
With this motivation, Chetrite et al. [2] consider a zero-mean Gaussian random velocity field process whose correlation matrix has entries of the form
(1.6) ${\mathbb{E}}\left[u_{i}(x,t)\,u_{j}(x^{\prime},t^{\prime})\right]\\ =\delta(t-t^{\prime})\sum_{{\mathbf{k}}\in\frac{2\pi}{P}{\mathbb{Z}}^{d}}\left[(1-\vartheta)\,\delta_{ij}-(1-\vartheta d)\,k_{i}k_{j}\right]e^{\text{\rm i}{\mathbf{k}}\cdot(x-x^{\prime})}\varrho(|{\mathbf{k}}|)$
for $1\leq i,j\leq d$. Here $\varrho$ is some spectral density function, and $0\leq\vartheta\leq 1$ is a parameter, called “degree of compressibility”, which equals zero if and only if the velocity field is divergence-free. The model is designed so that the spatial statistics of the resulting velocity field remain the same under every transformation of space that leaves the hypercube invariant; this has the remarkable consequence that, regardless of the dimension $d$, these statistics are characterised by just three parameters, denoted $\alpha$, $\beta$ and $\gamma$ by Chetrite et al., and constrained by the inequalities
$\gamma\geq|\beta|\;\;\text{and}\;\;4\alpha+(d+2)\beta+2\gamma\geq d|\beta|\,.$
The degree of compressibility may be expressed as
(1.7) $\vartheta=\frac{2\alpha+(d+1)\beta+2\gamma}{2\alpha+2\beta+d\gamma}$
so that, in the case where the velocity field is divergence free, only two “free” parameters remain. Chetrite et al. also define the “degree of anisotropy” of the random flow by
(1.8) $0,1]\ni\kappa:=\left|\frac{\alpha}{\alpha+\beta+\gamma}\right|\,.$
This terminology is justified by the fact that the spatial statistics of the velocity field are invariant under every rotation if and only if $\alpha=0$.
With this model, Equation (1.3) becomes a diffusion process on the group— $\text{SL}(d,{\mathbb{R}})$ if $\vartheta=0$, $\text{GL}(d,{\mathbb{R}})$ otherwise— that can be studied via its generator. This generator takes the form of a second-order partial differential operator in the variables that are used to parametrise the group. We shall return to Chetrite et al.’s analysis of this model in due course.
@@ 1.3. Renewing flows
In the present paper, we study an alternative model. We begin by considering a $d$-dimensional random velocity field with components of the form
@@
where
@@
    the $u_{ij}:{\mathbb{R}}\rightarrow{\mathbb{R}}$ are periodic functions with a common period $P/m$ for some $m\in{\mathbb{N}}$, and zero mean over a period;
@@
    the $\{\eta_{ij}(t)\}_{t\geq 0}$ are independent copies of the same random process $\{\eta(t)\}_{t>0}$;
@@
After selecting a suitable reference length, the common period of the $u_{ij}$ may be taken to equal $2\pi$ in dimensionless units. We shall henceforth assume that, for every $t>0$, the random variable $\eta(t)$ is uniformly distributed in $0,2\pi)$; this ensures that the resulting velocity field process is spatially homogeneous.
@@
(1.10) ${\mathcal{I}}_{n,i}:=\left(t_{n+1}-i\,\delta t,\,t_{n+1}-(i-1)\,\delta t\right)\,,\;\;i\in\{1,\,2,\,\ldots\,,d\}\,$,
the only non-vanishing velocity component is the $i$th one. Such flows are called “random flows with renewal” or, simply, “renewing flows” [3, 31].
As a concrete example, for $d=2$, one recovers the “sine flow” example considered by Haynes & Vanneste— see [16], Equations (3.17) and (4.8)— by taking both $u_{12}$ and $u_{21}$ to be some multiple of the sine function; the resulting equation for $\Gamma(x,t)$ is derived in their §4. Other examples may be found in [7].
@@ 1.4. Products of matrices from renewing flows
It is proved in Appendix A that the $\xi_{ij}$ are independent. We are thus led to studying certain products
@@
of random independent and identically-distributed matrices with a particular structure: Firstly, the $g_{n}$ have unit determinant; this is a consequence of the fact that the velocity field defined by Equation (1.9) is divergence-free. Secondly, because of the way we chose to discretise the dynamical system, each $g_{n}$ is itself a product of random independent triangular matrices; this will facilitate the analysis to follow.
@@ 1.5. The generalised Lyapunov exponent
Associated with the product (1.17) is the generalised Lyapunov exponent:
(1.18) $L(\ell):=\lim_{n\rightarrow\infty}\frac{1}{n}\ln{\mathbb{E}}\left[1_{\ell}\left(x\Pi_{n}\right)\right]=\sum_{j=1}^{\infty}c_{j}\frac{\ell^{j}}{j!}$
@@
$L$ may be viewed as a scaled cumulant generating function associated with the dimensionless random variable $\ln\left|{x}\Pi_{n}\right|$; it is in fact independent of the x{x} appearing in its definition. In terms of the renewing flow, $c_{j}/(\delta t\,d)$ is the discrete counterpart of
@@
so that, in particular, $c_{1}/(\delta t\,d)$ measures the rate at which nearby trajectories separate. More generally, $L$ encodes the large-$n$ statistical properties of the product: if $\text{\rm d}s$ denotes an interval of infinitesimal length $ds$ centered at $s\in{\mathbb{R}}$, then
(1.20) ${\mathbb{P}}\left(\frac{1}{n}\ln\left|x\Pi_{n}\right|\in\text{\rm d}s\right)\propto\exp\left[-n\,I(s)\right]\,ds\;\;\text{as $n\rightarrow\infty$}\,.$
where $I$, known variously as the “large deviation” [9] or “rate” [26] or “entropy” [8] or “Cramér” [16] function associated with the product, is the Legendre transform of $L$:
(1.21) $I(s):=\min_{\ell\in{\mathbb{R}}}\left[\ell s-L(\ell)\right]\,.$
@@ 1.6. The transfer operator
where $\lambda(\ell)$ is the dominant eigenvalue of the transfer operator. This characterisation of the generalised Lyapunov exponent assumes that there is a “spectral gap”— that is, that the dominant eigenvalue is well separated from the rest of the spectrum.
@@ 1.7. The continuum limit
For the piecewise-constant-in-time velocity field considered in this paper, only the diagonal entries of the velocity field’s correlation matrix are non-zero; a simple calculation shows that, for $1\leq i\leq d$, they take the form
(1.26) ${\mathbb{E}}\left[u_{i}(x,t)\,u_{i}(x^{\prime},t^{\prime})\right]\\ =\sum_{j\neq i}b_{ij}(x_{j}-x_{j}^{\prime})\,\begin{cases}1&\text{if $t,\,t^{\prime}\in{\mathcal{I}}_{n,i}$ for some $n\in{\mathbb{N}}$}\\ 0&\text{otherwise}\end{cases}$
where ${\mathcal{I}}_{n,i}$ is the time interval defined by Equation (1.10) and the $b_{ij}$ are cosine series with positive coefficients. If we multiply this identity by a suitable test function of $t^{\prime}$ and integrate over $t^{\prime}$, we see that the limit $\delta t\rightarrow 0+$, henceforth referred to as the continuum limit, corresponds to a velocity field that, as in the Kraichnan ensemble of §1.2, is delta-correlated in time.
@@ 1.8. Outline of the paper and main results
Equation (1.25) then implies that
[…]
@@
(1.34) $\Lambda(k,\ell)=\sum_{j=1}^{\infty}\gamma_{j}(k)\frac{\ell^{j}}{j!}:=(d-1)\,\ell+\frac{d-1}{d}\,\ell^{2}+\mu(k,\ell)$
and $\mu(k,\ell)$ is the particular eigenvalue that maximises the expression on the right-hand side. We denote by
(1.35) $\Upsilon(k,s):=\min_{\ell\in{\mathbb{R}}}\left[\ell s-\Lambda(k,\ell)\right]$
the Legendre transform of $\Lambda(k,\ell)$. We then take advantage of the fact that the companion spectral problem (1.32) is completely solvable when $k=0$ by embedding our original problem in a parameter-dependent family, where the parameter $k$ assumes “small” values. To facilitate the calculations, we work with the Iwasawa realisation, described in §3, of the basic representation, because it is in this realisation that the eigenfunctions of the Laplacian assume their simplest form.
In §4, we study the $k$-dependent spectral problem (1.32) when $d=2$. First, for fixed $0\leq k<1$, we consider its small-$\ell$ limit. By making a change of variable, we obtain a problem that is solvable when $\ell=0$ and this eventually leads to explicit formulae for the first two cumulants:
(1.36) $\gamma_{1}(k)=2\,\frac{\boldsymbol{\mathsf{E}}}{\boldsymbol{\mathsf{K}}}-1$
@@
(1.37) $\frac{\gamma_{2}(k)}{2}=\frac{3}{2}-k^{2}-\frac{\boldsymbol{\mathsf{E}}}{\boldsymbol{\mathsf{K}}}+\frac{2\pi^{2}}{\boldsymbol{\mathsf{K}}^{2}}\sum_{j=1}^{\infty}\left(\frac{q^{2j}}{1-q^{4j}}\right)^{2}-\frac{6\pi^{2}}{\boldsymbol{\mathsf{K}}^{2}}\sum_{j=1}^{\infty}\left(\frac{q^{2j-1}}{1-q^{4j-2}}\right)^{2}$
where $\boldsymbol{\mathsf{K}}$ and $\boldsymbol{\mathsf{E}}$ are the complete elliptic integrals of the first and second kind respectively, evaluated at $k$, and $q$ is the corresponding nome; see Equation (4.14).
Equations (1.36-1.37) essentially reproduce formulae obtained previously by Chetrite et al. for the two-dimensional Kraichnan model; see their sections 4.4 and 4.5 [2]. Interestingly, Formula (1.36) also appears in connection with Anderson’s model at zero energy and the Dirac equation with a random mass [10, 24, 23]. On the purely mathematical level, the study of the large-deviation statistics in all these models reduces to the same spectral problem (1.32); as we shall see, the only difference is that, in our model, the parameter $k$ assumes real rather than imaginary values.
The fact that the parameter $k$ appears in formulae (1.36-1.37) as the modulus of the elliptic integrals gives some confidence that its introduction may indeed be the “right” way of relating our spectral problem to that for $\Delta_{K}$. Plots of
(1.38) $\ell\mapsto\ell\,\gamma_{1}\left(k\right)+\ell^{2}\,\frac{\gamma_{2}\left(k\right)}{2}=\Lambda\left(k,\ell\right)+o\left(\ell^{2}\right)$
@@
(1.39) $s\mapsto\frac{\left[s-\gamma_{1}\left(k\right)\right]^{2}}{2\gamma_{2}\left(k\right)}=\Upsilon\left(k,s\right)+o\left(s^{2}\right)$
against their respective argument, for the case $k=1/\sqrt{2}$, are shown in Figure 1 (a). These formulae express the fact that, in the continuum limit, the distribution of the random variable $\left|x\Pi_{n}\right|$ is approximately lognormal for large $n$ [8].
Refer to caption(a)(b)
Figure 1. Plots of $\Lambda(1/\sqrt{d},\ell)$ against $\ell$ (solid blue curve) and $\Upsilon(1/\sqrt{d},s)$ against $s$ (solid red curve) for (a) $d=2$ and (b) $d=3$. In (a) are also displayed the approximation (1.38) of $\Lambda(1/\sqrt{2},\ell)$ (dashed curve) and the approximation (1.39) of $\Upsilon(1/\sqrt{2},s)$ (dot-dashed curve) which use only the first two cumulants.
Then, by considering the small-$k^{2}$ limit of the spectral problem (1.32), we obtain expansions for $\Lambda$ and $\Upsilon$ in ascending powers of $k^{2}$:
(1.40) $\Lambda(k,\ell)=\frac{\ell\left(\ell+2\right)}{2}-\frac{\ell\left(\ell+2\right)}{2}k^{2}\\ +\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{128}k^{4}+\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{256}k^{6}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(7\ell^{4}+28\ell^{3}+36\ell^{2}+16\ell-5248\right)}{2097152}k^{8}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(21\ell^{4}+84\ell^{3}+108\ell^{2}+48\ell-7552\right)}{4194304}k^{10}+\cdots$
@@
(1.41) $\Upsilon(k,s)=\frac{\left(s-1\right)^{2}}{2}+\frac{\left(s-1\right)\left(s+1\right)}{2}k^{2}\\ -\frac{s^{4}-74s^{2}+9}{128}k^{4}-\frac{\left(3s^{2}-14s+3\right)\left(3s^{2}+14s+3\right)}{256}k^{6}\\ +\frac{7s^{8}+948s^{6}-211974s^{4}+1781972s^{2}-47241}{2097152}k^{8}\\ +\frac{133s^{8}+15244s^{6}-990274s^{4}+4389324s^{2}-67995}{4194304}k^{10}+\cdots$
In §5, we turn to the case $d=3$. In this case, there is no obvious change of variables that makes the problem tractable when $\ell=0$, and so we are unable to make progress in the small-$\ell$ limit. In the small-$k^{2}$ limit, however, we can once again compute expansions for $\Lambda$ and its Legendre transform $\Upsilon$; we obtain
(1.42) $\Lambda(k,\ell)=\frac{2\ell\left(\ell+3\right)}{3}-\frac{4\ell\left(\ell+3\right)}{5}k^{2}\\ +\frac{12\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)}{875}k^{4}\\ -\frac{72\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)\left(9\ell^{2}+27\ell-395\right)}{3128125}k^{6}+\cdots$